# CachePrune: Neural-Based Attribution Defense Against Indirect Prompt Injection Attacks

### v1

@@ Abstract
Large Language Models (LLMs) are identified as being susceptible to indirect prompt injection attack, where the model undesirably deviates from user-provided instructions by executing tasks injected in the prompt context. This vulnerability stems from LLMs’ inability to distinguish between data and instructions within a prompt. In this paper, we propose CachePrune that defends against this attack by identifying and pruning task-triggering neurons from the KV cache of the input prompt context. By pruning such neurons, we encourage the LLM to treat the text spans of input prompt context as only pure data, instead of any indicator of instruction following. These neurons are identified via feature attribution with a loss function induced from an upperbound of the Direct Preference Optimization (DPO) objective. We show that such a loss function enables effective feature attribution with only a few samples. We further improve on the quality of feature attribution, by exploiting an observed triggering effect in instruction following. Our approach does not impose any formatting on the original prompt or introduce extra test-time LLM calls. Experiments show that CachePrune significantly reduces attack success rates without compromising the response quality. Note: This paper aims to defend against indirect prompt injection attacks, with the goal of developing more secure and robust AI systems.
@@ 1 Introduction
Figure 1: Illustration of indirect prompt injection attack with LLMs.
@@
The rapid advancements in Large Language Models (LLMs) Achiam et al. (2023); Touvron et al. (2023) have revolutionized natural language processing (NLP) for a wide range of tasks Becker et al. (2024); Upadhyay et al. (2024). However, these models exhibit a critical vulnerability on indirect prompt injection attacks Yi et al. (2023); Greshake et al. (2023), where instructions injected within the prompt context can override user-provided directives (Figure 1). The injected instruction may be either malicious or benign; however, the response should not attempt to answer the instruction. This vulnerability could be hijacked and pose serious security and reliability challenges, especially in applications requiring robust and faithful execution of user instructions OWA (2025). Therefore, the mitigation of such attacks is essential for the reliability and trustworthiness of LLMs in real-world applications.
@@
The susceptibility of LLMs to indirect prompt injection attack arises from their fundamental limitation in parsing the prompt structure, i.e., unable to distinguish between data and instructions within a prompt Zverev et al. (2024); Chen et al. (2024a). In defending against such attacks, re-training the LLMs Chen et al. (2024a); Piet et al. (2024); Chen et al. (2024b) to adhere to the prompt structure can be computationally prohibitive. Alternatively, existing mitigation strategies focus on imposing rigid prompt formatting with reminder instructions (Wu et al., 2023; Hines et al., 2024) or implementing supplementary test-time workflows (Wang et al., 2024; Jia et al., 2024), so that the user requests are prioritized in response generation. Such modifications often result in limited defense effects or incurring supplementary test-time computation with extra LLM calls for each response processed. Additionally, they could potentially interfere with the intended user instructions, thereby undermining the quality of the model’s output. These limitations highlight the need for alternative and efficient solutions that mitigate the attacks while not overly altering the original prompt and its response generation workflow.
In this paper, we focus on the source of LLMs’ vulnerability, i.e., the model’s confusion between user-specified context (data) and instruction. We start our approach with a fundamental question: What makes the difference between data and instruction from the LLMs’ perspective? Different from the user, the LLM has its own way of telling between data and instructions. Specifically, a text span is identified as an instruction by the LLM if the model responds to it by giving a solution, while it is identified as data if the model only leverages its content as supportive information. The indirect prompt injection attack occurs when such definition misaligns with the user-defined spans of context and instruction. To solve this misalignment, we propose CachePrune that 1) identifies neurons that can make a difference between data and instruction for the LLM, and 2) prune to enforce such difference between the user-defined context and instruction, so the context span is only treated as supportive information instead of any indicator of instruction following. Specifically, we leverage feature attribution for 1) that attributes the model generations back to the neurons in the Key-Value (KV) cache of the prompt context. Notably, our analysis reveals that the execution of injected instructions relies on the activating of only a small subset of neurons (e.g., 0.5%). For 2), we prune such neurons on the span of input context from the prompt KV cache. In this way, we enforce the LLM to interprete the input context exclusively as pure data, thus mitigating the risk of responding to its injected instructions.
@@
Our feature attribution relies on a proposed loss function induced from an upperbound of the Direct Preference Optimization (DPO) Rafailov et al. (2024) objective. We show that it is sample efficiency that enables effective feature attribution with only a few samples. We further improve on the quality of feature attribution leveraging an observed trigerring effect in the instruction following. Notably, our proposed CachePrune is lightweight, requiring only a mask of identified neurons for pruning, without introducing extra test-time computation or LLM calls per response. It is complementary to the existing defensive approaches that modify the original prompt or workflow of response generation.
@@
  • We propose CachePrune that mitigates indirect prompt injection attack, by identifying and pruning neurons in the context KV cache that trigger instruction following. This enforces the LLM to treat the input context as pure data, thus preventing the LLM from responding to the injected instruction.
  • In identifying these neurons, we propose a feature attribution mechanism with a loss function that enables effective attribution with only few samples. We also leverage an observed triggering mechanism that further improves the quality of feature attribution.
@@ 2.1 Preliminary
Prompting LLMs: Let $x=[x_{t}]_{t=1}^{T}\sim\mathcal{X}$ be an input prompt with $T$ tokens, consisting of the user-specified instruction and its context. $p_{\theta}(\cdot|x)$ is the output probability with an LLM of $L$ layers parameterized by $\theta$. The LLM is expected to answer the user-specified instruction, leveraging the context as data that provides supporting information. Here, the data may not necessarily be numeric, but a text span of supporting information that helps solve the instruction.
State-of-the-art LLMs generally adopt the Transformer Waswani et al. (2017) architecture, where each token $x_{t}$ is encoded by layer $l$ into a key vector $k_{t,l}\in\mathcal{R}^{D}$ and a value vector $v_{t,l}\in\mathcal{R}^{D}$. Let $\mathcal{H}_{x}=[h_{t}]_{t=1}^{T}$ be the KV cache of prompt $x$, where $h_{t}=[k_{t,1};v_{t,1};\cdots;k_{t,L};v_{t,L}]\in\mathcal{R}^{2\times D\times L}$ is the concatenation of key and value vectors from all layers in step $t$. For a length-$K$ response $y=[y_{t}]_{t=1}^{K}\in\mathcal{Y}$, $y_{t}$ is generated with,
@@
where $y_{<t}$ denotes the response tokens up to step $t$. $\mathcal{H}_{x}$ is independent of the response, and thus is reused with different $y_{t}$.
Indirect Prompt Injection Attack: In an indirect prompt injection attack, the prompt context is injected with instructions from third party. The injected instructions can be malicious or not, but are not intended to be responded by the LLM. As illustrated in Figure 1, we define $y^{p}\sim\mathcal{Y}^{p}_{x}$ as a poisoned response of $x$, if $y_{p}$ deviates from the user instruction but responds to the injected ones from context. Similarly, we define $y^{c}\sim\mathcal{Y}^{c}_{x}$ as a clean response of $x$ if $y^{c}$ ignores the injected instructions. In evaluation, we call an LLM being subjected to indirect prompt injection attack with $x$ if,
@@
||$\cdot$|| ⋅ | is the support of a distribution. $y^{*}$ can be approximated with greedy sampling.
Defending against an indirect prompt injection attack can be characterized as promoting $y_{c}$ over $y_{p}$ in the response generation. In this paper, we achieve it by identifying and pruning neurons of the KV Cache that trigger the LLM responding to the injected instructions in context. Our approach that prunes on the KV Cache is compatible with context caching gem (2025); ope (2025), e.g., enabling efficient prompting when there are multiple questions/instructions with the same cached context. For such cases, the KV-Cache of the context or prompts only needs to be pruned once, then reliably saved for future LLM calls without worrying about being attacked by its injections.
@@ 2.2 Defending Against Indirect Prompt Injection Attack
Figure 2: Illustration of our workflow of defending against indirect prompt injection.
To defend against indirect prompt injection attacks, our approach leverages the LLM’s inherent ability to distinguish between data and instructions, enforcing this distinction so the LLM can recognize the boundary between user-provided context and instruction.
Specifically, as discussed in Section 1, an input text span is considered data if the LLM leverages it solely as supporting information without following it as an instruction. Therefore, we characterize the distinction between data and instruction as features in the prompt KV cache that trigger the LLM to react on injected instructions. We identify these features through feature attribution and selectively prune them from the input context span, ensuring that the LLM interprets the context as pure data. By enforcing the LLM’s awareness on the boundary between input context and instructions, this approach enhances the model’s resilience against prompt injection attacks.
Feature Attribution. We aim at identifying key features within the prompt’s KV cache that contribute to the observed difference in the LLM’s behavior, i.e., when interpreting the input context as either pure data or as instructions. Let $\mathcal{L}^{attr}:\mathcal{Y}^{c}_{x}\times\mathcal{Y}^{p}_{x}\times\mathcal{X}\rightarrow\mathcal{R}$ be an attribution loss function that captures such difference in LLM outputs. Specifically, we have a larger $\mathcal{L}^{attr}$ indicating the LLM mistakenly treats the input context as instructions, i.e., by preferring a poisoned response $y_{p}$ over the clean one $y_{c}$. For clarity, we defer the details of the loss function to Section 2.3.
In attributing $\mathcal{L}^{attr}$ to features in the KV cache, we follow Shrikumar et al. (2017); Yang et al. (2022) that score each feature by its contribution to $\mathcal{L}^{attr}(\mathcal{Y}^{c}_{x},\mathcal{Y}^{p}_{x};\mathcal{X})$. Let $h^{i}_{t}$ be the $i$th feature of the key-value vector $h_{t}$, which is scored by,
@@
In our approach, we perform feature attribution solely on the input context span, as injected instructions are embedded exclusively within the input context. Let $c_{s}$ and $c_{e}$ be the input token index that marks the start and end of user-specified context, respectively. We denote $\mathcal{A}=[a_{t}]_{t=c_{s}}^{c_{e}}$ be our attribution matrix and $a_{t}=[a_{t}^{i}]_{i=1}^{2\times D\times L}$ is the attribution vector for $h_{t}$. In the experiments, we compute $\mathcal{A}$ with $N=8$ samples. $h_{t>{c_{e}}}$ will be generated with the pruned $\mathcal{[}h_{t}]_{t=c_{s}}^{c_{e}}$ during testing.
Aggregation. Note that $h^{i}_{t}$ with the same dimension $i$ originate from the same neuron. Thus, we aggregate attribution scores for each neuron by taking the maximum value across time steps,
@@
A neuron with large $a^{i,neu}$ is more influential on the LLM’s output, i.e., either treating the context as pure data or reacting on its injected instructions. This answers the question raised in Section 1: The activation of such neurons makes a difference on the LLM’s view over data vs instruction, by triggering the LLM to respond to the injected instruction in context.
@@
[…]
$\tau=\text{sup}_{\tau\in\mathbb{R}}Pr(a^{i,neu}\geq\tau,i\in\Phi)\,\,\ \leq\,\,p$ (5)
$Pr$ is uniform over all neurons. Each dimension $i$ is masked with $m_{i}$ that,
@@
where $\mathbbm{1}$ is the indicator function with $\alpha$ defaults as 1. $m_{i}$ reflects the LLM’s own recognition of what differentiates pure data from instructions. Applying this mask over the context enforce the LLM interprete the context solely as data, thus mitigating indirect prompt injection attack. We abstract our workflow in Figure 2.
2.3 The Attribution Loss
According to Section 2.2, the feature attribution is guided by an attribution loss $\mathcal{L}^{attr}$, which quantifies the observed difference between interpreting the input context as either pure data or instructions. This can be evaluated as an objective of preference optimization, among which the most common and effective one is the Direct Preference Optimization (DPO) Rafailov et al. (2024).In the context of indirect prompt injection, the DPO objective $\mathcal{L}_{DPO}$ can be defined as,
@@ 2.3 The Attribution Loss
where $\beta>0$, $\mathcal{D}=\{\mathcal{X},\mathcal{Y}^{c}_{x},\mathcal{Y}^{p}_{x}\}$ is the perference optimization dataset. $\sigma(\cdot)$ is the sigmoid function. An accurate estimation of $\mathcal{L}_{DPO}$ will ideally capture the difference between the LLM’s perception on data vs. instruction. Specifically, a higher $\mathcal{L}_{DPO}$ indicates the context being mistakenly perceived as instruction, and vice-versa.
However, the LLM’s output complexity grows exponentially with the response length. As a result, it requires substantial sampling and computation to estimate the expectation in (7) with low variance. In this paper, we instead derive our loss function from an upperbound of $\mathcal{L}_{DPO}$.
where ||$\cdot$|| ⋅ | is the support of a distribution. $\mathcal{C}_{ref,\mathcal{D}}$ is a constant to $\theta$ that is functioned by $\mathcal{D}$ and the DPO reference model $ref$. $\mathbb{H}(\mathcal{Y}^{c}_{x}|x)$ and $\mathbb{H}(\mathcal{Y}^{p}_{x}|x)$ are the entropy of clean and poisoned responses given $x$. $p_{\theta}(y\in|\mathcal{Y}^{p}_{x}|\,|x)$ is the gross probability of generating poisoned responses from $x$, and similar to $p_{\theta}(y\in|\mathcal{Y}^{c}_{x}|\,|x)$. The proof is in Appendix A.
Especially, the clean and poison probabilities $p_{\theta}(y\in|\mathcal{Y}^{p/c}_{x}|\,|x)$ are not sufficient to capture the objective of preference optimization. In order to minimize (8), we should also attend to the uniformity with $\mathbb{H}(\mathcal{Y}^{p/c}_{x}|x)$. This requires computing the expectation over the generated responses,which is sample inefficient due to the complexity of the space of generated responses. Here, we delegate the entropy terms with the most probable poison and clean responses, denoted as $y^{p/c,*}_{x}=\text{argmax}_{y\in|\mathcal{Y}_{x}^{p/c}|}p_{\theta}(y|x)$. Intuitively, given the gross probability $p_{\theta}(y\in|\mathcal{Y}^{p/c}_{x}|\,|x)$, $\mathbb{H}(\mathcal{Y}^{p/c}_{x}|x)$ should be generally lowered if $y^{p/c,*}_{x}$ gets higher probability, vice versa. Therefore, we define an attribution loss that is inspired from (8), i.e.,
@@
where full denotes feature attribution with all response tokens, which will be discussed later.
We can observe that (9) captures both the objectives of probability and uniformity in (8): a) Minimizing (9) promotes $p_{\theta}(y\in|\mathcal{Y}^{c}_{x}|\,|x)$, while suppressing $p_{\theta}(y\in|\mathcal{Y}^{p}_{x}|\,|x)$. b) Given the gross probability $p_{\theta}(y\in|\mathcal{Y}^{p/c}_{x}|\,|x)$, we have
Lemma 1. $\mathcal{L}^{attr}_{full}$ that is ranged between $[-1,1]$ is closely associated with $\mathcal{L}_{DPO}^{u}$ by,
We can obseve that Lemma 1 will no longer hold with only few samples, if we follow (7) that replace $y^{p/c,*}$ with $y^{p/c}$ in $\mathcal{L}_{full}^{attr}$. This suggests to sample with the most probable responses for feature attribution. In experiemnt, we sample $y^{p,*}$ with greedy decoding when $p_{\theta}(y^{p,*}|x)>p_{\theta}(y^{c,*}|x)$, then $p_{\theta}(y^{c,*}|x)$ is approximated also using greegy decoding but with injected instruction removed. Conversely, we sample $y^{c,*}$ with greedy decoding when $p_{\theta}(y^{p,*}|x)<=p_{\theta}(y^{c,*}|x)$. In this case, $y^{p,*}$ is approximated by concatenating the injected instruction with user queries, so that the model cannot ignore the injected instruction. Please refer to Figure 6 for details.
The subset $\Phi$. With (9), we can find that the attribution score (3) can be decomposed into,
@@
$a_{t,p}^{i}$ and $a_{t,c}^{i}$ are the scores for poisoned and clean contributions. Correspondinglly, we can have $a_{p}^{i,neu}=max_{t}\,\,a_{t,p}^{i}$ and $a_{p}^{i,neu}=max_{t}\,\,a_{t,c}^{i}$. We want to avoid pruning on neurons with significant clean contribution, so that the pruned LLM can generate clean responses that address the user instruction. Let $a_{p}^{i,norm}=a_{p}^{i,neu}/\sum_{i^{\prime}}a_{p}^{i^{\prime},neu}$ and $a_{c}^{i,norm}=a_{c}^{i,neu}/\sum_{i^{\prime}}a_{c}^{i^{\prime},neu}$ be the normalized contribution scores, we have $\Phi$ defined by,
@@
Figure 3: The rank of predicted response tokens. Taking (a) as an example, "clean greedy" means the response from greedy decoding is clean. Therefore, the tokens from $y^{*}_{c}$ are always ranked zero. In this case, a clean response can be triggered with only one or two tokens. Note that we assume $y_{x}^{p,*}$ starts with answering the injected instructions. We do not find such effect when the injected instructions are answer in the end.

### v2

@@ Abstract
Large Language Models (LLMs) are susceptible to indirect prompt injection attack, in which the model inadvertently responds to task messages injected within the prompt context. This vulnerability stems from LLMs’ inability to distinguish between data and instructions within a prompt. In this paper, we propose CachePrune that defends against this attack by identifying and pruning task-triggering neurons from the KV cache of the input prompt context. By pruning such neurons, we encourage the LLM to interpret the input prompt context purely as data rather than any cues for instruction following. To identify these neurons, we introduce a neural attribution mechanism guided by a preferential attribution loss, which enables effective attribution with only a few samples while preserving response quality after pruning. We further enhance the efficacy of neural attribution by leveraging an observed triggering effect inherent in the model’s response generation behavior. Notably, our approach does not impose additional formatting on the prompt or introduce extra test-time LLM calls. Experiments show that CachePrune can significantly reduce attack success rates while not compromising the response quality.
CachePrune: Neural-Based Attribution Defense Against Indirect Prompt Injection Attacks
Rui Wang1   Junda Wu2   Yu Xia2   Tong Yu1  Ruiyi Zhang1   Ryan Rossi1 Subrata Mitra1   Lina Yao3,4   Julian McAuley2 1Adobe Research 2University of California San Diego 3University of New South Wales  4CSIRO’s Data61 {ruiwan,tyu,ruizhang,ryrossi,sumitra}@adobe.com {juw069,yux078,jmcauley}@ucsd.edu, lina.yao@unsw.edu.au
@@ 1 Introduction
Figure 1: Illustration of indirect prompt injection attack with LLMs. The attack message is shown in orange.
@@
The rapid advancements in Large Language Models (LLMs) achiam2023gpt; touvron2023llama have revolutionized natural language processing (NLP) for a wide range of tasks becker2024text; upadhyay2024comprehensive. However, existing LLMs exhibit critical vulnerability to indirect prompt injection attacks yi2023benchmarking; greshake2023not; abdelnabi2024you, where instructions or task messages injected within the prompt context can override user-provided directives (Figure 1). This vulnerability can be exploited by malicious actors, posing significant risks to the reliability and trustworthiness of LLMs in real-world applications OWASP.
@@
The susceptibility of LLMs to indirect prompt injection attacks stems from their fundamental limitation in parsing prompt structure, i.e., their inability to distinguish between data and instructions within a prompt zverev2024can; chen2024struq. In defending against such attacks, re-training the LLMs chen2024struq; piet2024jatmo; chen2024aligning to adhere to the prompt structure can be computationally prohibitive. Alternatively, existing mitigation strategies focus on imposing rigid prompt formatting with reminder instructions (wu2023defending; hines2024defending) or implementing supplementary test-time workflows (wang2024fath; jia2024task), so that the user instructions can be prioritized in response generation. Such modifications often result in limited defense efficacy or incurring test-time computation overheads that require extra LLM calls for each response processed. Additionally, they could potentially interfere with the users’ intended instructions, thereby undermining the quality of the model’s output. These limitations underscore the need for alternative and efficient solutions that mitigate such attacks while not overly altering the original prompt and its response generation workflow.
In this paper, we focus on the source of LLMs’ vulnerability, i.e., the model’s confusion between context (data) and instructions specified by the prompt structure. We start our approach with a fundamental question: What makes the difference between data and instructions from the LLMs’ perspective? Different from the user, the LLM has its own way to differentiate between data and instructions. From a behavioral perspective, a text span is interpreted as an instruction if the model generates a response to it, and as data if its content is merely used as supportive information. An indirect prompt injection attack arises when the LLM-based differentiation misaligns with the user-defined separation between context and instructions. Our approach is motivated by solving such a misalignment with the following two general steps,
  • Neural Attribution: Identify neurons whose activations account for the LLM-based differentiation between context and instructions. These neurons trigger instruction following on the context, capturing the LLM’s internal insights on context vs. instructions.
  • Intervention: Intervene by pruning the identified neurons over the prompt context. It promotes alignment by enforcing the LLM-based differentiation between user-defined spans of instruction and context, ensuring that the context is treated purely as supportive information rather than cues for instruction following.
Since we identify neurons from the KV cache of the prompt context, our approach is dubbed CachePrune. Such an approach leaves the prompt formatting unchanged and is computationally lightweight, relying solely on a mask for pruning without incurring extra test-time computation per response or LLM calls. However, the challenges are twofold: a) How to define an attribution loss that effectively captures the LLM’s insights on context vs. instructions? b) How to regularize the attribution so that the pruning does not impair response quality on user instructions?
@@
In our approach, we introduce a preferential attribution loss that shares insights with an upperbound of the Direct Preference Optimization (DPO) rafailov2024direct objective. We show that it is sample efficient that enables effective attribution with only a few samples. Building on this, we impose a regularization on neuron attribution based on gradients from different loss terms, so that the pruning preserves response quality. We further improve on the quality of neural attribution leveraging an observed trigerring effect in response generation. In particular, since our CachePrune does not modify the prompt or inference workflow, it is complementary to the existing defensive approaches that require modifying the prompt formatting or workflows of response generation.
@@
  • We propose CachePrune that mitigates indirect prompt injection attack, by identifying and pruning neurons in the context KV cache that trigger instruction following. This enforces the LLM to treat the input context as pure data, thus preventing the LLM from responding to the injected instructions.
  • In identifying these neurons, we propose a neuron attribution mechanism with a loss function that enables effective attribution using only few samples, while preserving the response quality. We also leverage an observed triggering mechanism that further improves the quality of neural attribution.
@@ 2.1 Preliminary
Prompting LLMs: Let $x=[x_{t}]_{t=1}^{T}\sim\mathcal{X}$ denote an input prompt with $T$ tokens, consisting of the user-specified instruction and its context. $p_{\theta}(\cdot|x)$ is the output probability with an LLM of $L$ layers parameterized by $\theta$. The model is expected to answer the user-specified instruction, while leveraging the context as auxiliary data that provides supporting information.
State-of-the-art LLMs generally adopt the Transformer waswani2017attention architecture, where each token $x_{t}$ is encoded by layer $l$ into a key vector $k_{t,l}\in\mathcal{R}^{D}$ and a value vector $v_{t,l}\in\mathcal{R}^{D}$. Let $\mathcal{H}_{x}=[h_{t}]_{t=1}^{T}$ be the KV cache of prompt $x$, where $h_{t}=[k_{t,1};v_{t,1};\cdots;k_{t,L};v_{t,L}]\in\mathcal{R}^{2\times D\times L}$ is the concatenation of key and value vectors from all layers in step $t$. For a length-$K$ response $y=[y_{t}]_{t=1}^{K}\in\mathcal{Y}$, $y_{t}$ is generated with,
@@
where $y_{<t}$ denotes the response tokens up to step $t$. $\mathcal{H}_{x}$ is reused during inference with different $y_{t}$.
Indirect Prompt Injection Attack: In an indirect prompt injection attack, the injected instructions, whether malicious or benign, are not meant to elicit a response from the LLM. As illustrated in Figure 1, we define $y^{p}\sim\mathcal{Y}^{p}_{x}$ as a poisoned response of $x$, if it diverges from the user-specified instruction and follows the injected instructions present in the context. Similarly, we define $y^{c}\sim\mathcal{Y}^{c}_{x}$ as a clean response of $x$, if $y^{c}$ ignores the injected instructions. During evaluation, an LLM is considered attacked with $x$ if,
@@
$|\cdot|$ is the support of a distribution. $y^{*}$ can be approximated with greedy sampling.
Defending against an indirect prompt injection attack can be characterized as promoting $y_{c}$ over $y_{p}$ in the response generation. In this paper, we achieve it by identifying and pruning neurons of the KV Cache that trigger the LLM responding to the injected instructions. Importantly, our approach that prunes on the KV Cache is compatible with context caching gemin-cc; openai-cc, e.g., enabling efficient prompting when there are multiple questions/instructions with the same cached context. For such cases, the KV-Cache of the context or prompts only needs to be pruned once, then reliably saved for future LLM calls without worrying about being attacked by its injections.
@@ 2.2 Defending Against Indirect Prompt Injection Attack
Figure 2: Illustration of our workflow of CachePrune. Given a prompt with injection, our pruning is guided by a preferential attribution loss computed from sampled clean and poisoned responses. Then, we conduct neural attribution using the cached activations from forward-propagation and their gradients from the preferential attribution loss ( Blue). During intervention, a neuron is pruned by masking its corresponding row in the context KV cache ( Red). Activations after step $c_{s}$ should be generated on the pruned cache. $\lor$ means the preferential attribution loss is larger when the poisoned response is preferred.
Our CachePrune captures the LLM’s inherent differentiation between data and instructions, by identifying neurons that trigger the LLM’s response to injected instructions. We enforcing such differentiation between user-provided context and instruction, so that the context is only interpreted as supportive information instead of any clues for instruction following.
The framework is illustrated in Figure 2, where we identify these neurons by learning a pruning mask via neural attribution, then we intervene by applying the learnt pruning mask over the KV cache of the prompt context.
2.2.1 Neural Attribution
We aim at identifying neurons that contribute to the LLM-based differentiation on context and instructions. The identified neurons are included in a pruning mask derived through the following three steps. Notable, we regularize with Selective Thresholding for the preservation of response quality.
Attribution Scoring. Let $\mathcal{L}^{attr}:\mathcal{Y}^{c}_{x}\times\mathcal{Y}^{p}_{x}\times\mathcal{X}\rightarrow\mathcal{R}$ be an attribution loss function that captures such difference in LLM outputs. Specifically, we have a larger $\mathcal{L}^{attr}$ indicating the LLM mistakenly treats the input context as instructions, i.e., by preferring a poisoned response $y_{p}$ over the clean one $y_{c}$. For clarity, we defer the details of the loss function to Section 2.3.
In attributing $\mathcal{L}^{attr}$ to neurons in the KV cache, we follow shrikumar2017learning; yang2022task that score each neuron activation by its contribution to $\mathcal{L}^{attr}(\mathcal{Y}^{c}_{x},\mathcal{Y}^{p}_{x};\mathcal{X})$. Let $h^{i}_{t}$ be the $i$th neuron activation in the key-value vector $h_{t}$, which is scored by,
@@ 2.2.1 Neural Attribution
In our approach, we perform feature attribution solely on the input context span, as injected instructions are embedded exclusively within the input context. Let $c_{s}$ and $c_{e}$ be the input token index that marks the start and end of user-specified context, respectively. We denote $\mathcal{A}=[a_{t}]_{t=c_{s}}^{c_{e}}$ as our attribution matrix and $a_{t}=[a_{t}^{i}]_{i=1}^{2\times D\times L}$ is the attribution vector for $h_{t}$. In the experiments, we compute $\mathcal{A}$ with $N=8$ samples. $h_{t>{c_{e}}}$ will be generated with the pruned $\mathcal{[}h_{t}]_{t=c_{s}}^{c_{e}}$ during testing.
Neural Aggregation. Note that all $h^{i}_{t}$ corresponding to the same neuron share the same $i$. Thus, we aggregate attribution scores for each neuron by taking the maximum value across time steps,
@@
Selective Thresholding. Intervening on neurons with large $a^{i,neu}$ should alter the LLM’s interpretation between data and instruction. However, this would degrade the response quality since our $\mathcal{L}^{attr}$ is only about poison vs. clean, without regularizion for quality preservation.
@@
$\tau=\text{inf}_{\tau_{0}\in\mathbb{R}}\,\mathbb{E}_{i}\,(\mathbbm{1}\{a^{i,neu}\geq\tau_{0},i\in\Phi\})\ \leq\,\,p$ (5)
where $\mathbbm{1}$ is the indicator function. For the $i$th neuron, its masking value is $\mathbbm{1}\{a^{i,neu}\geq\tau,i\in\Phi\})$.
2.2.2 Intervention
In experiments, the pruning mask is learnt from neural attribution with only few samples, then applied to all the prompts during testing. Specifically, we enforce the boundary between the user-defined context and instructions by applying the pruning mask over the prompt context. For each $h_{t}^{i}$ from the context KV cache with $t\in[c_{e},c_{s}]$, we apply the masking by multiplying with $m_{i}$,
@@ 2.2.2 Intervention
where $\alpha$ defaults to 1. $m_{i}$ reflects the LLM’s own recognition of what differentiates data from instructions. Applying this mask over the context enforce the LLM interprete the context solely as data, thus mitigating indirect prompt injection attack.
2.3 The Preferential Attribution Loss
According to Section 2.2, the neural attribution is guided by a preferential attribution loss $\mathcal{L}^{attr}$, which quantifies the observed difference between interpreting the context as either pure data or instructions. This can be evaluated as an objective of preference optimization, among which the most common and effective one is the Direct Preference Optimization (DPO) rafailov2024direct. In the context of indirect prompt injection, the DPO objective $\mathcal{L}_{DPO}$ can be defined as,
@@ 2.3 The Preferential Attribution Loss
where $\beta>0$, $\mathcal{D}=\{\mathcal{X},\mathcal{Y}^{c}_{x},\mathcal{Y}^{p}_{x}\}$ is the perference optimization dataset and $p_{ref}$ is a reference model. $\sigma(\cdot)$ is the sigmoid function. An accurate estimation of $\mathcal{L}_{DPO}$ will ideally capture the difference between the LLM’s perception on data vs. instruction. Specifically, a higher $\mathcal{L}_{DPO}$ indicates the context being mistakenly perceived as instruction, and vice-versa.
Our preferential attribution loss is a simplified variant of (7). Let $y^{p,*}_{x}$ and $y^{c,*}_{x}$ be the most probable poisoned and clean responses from prompot $x$, respectively. We define our preferential attribution loss as,
@@
full denotes attributing with full response, which will be discussed in Section 2.4. Specifically, we employ the following practical simplification:
  • We directly compute feature attribution using the prediction probability without logarithm. Note that this is consistent with the previous work, e.g., yang2023mitigating, that computes feature attribution with the predicted probability instead of the loglikelihood. [^1] We suspect that this is because of the logarithm distorts the allocation of attributed scores, overemphasizing a few prominent features that may include some false positives. We believe it is an interesting direction for future works. [^2] No $p_{ref}$ since our goal is to identify neuron contributions, i.e., not to regularize the extent of parameter updates.
  • We leverage the most probable $y^{p,*}_{x}$ and $y^{c,*}_{x}$ instead of random samples for sample efficiency, as the most probable responses better represent the output distributions. We further analyze in Appendix B by deriving an upperbound of (7) in Theorem 1. Together with Lemma 1, we show that (8) with $y^{p,*}_{x}$ and $y^{c,*}_{x}$ is reasonable, and is consistent with the upperbound in the asymptotic cases.
The most probable clean and poisoned responses cannot be sampled exactly. We approximate by following the greedy positive/negative sampling in Figure 7, which is further detailed in Appendix E.
The subset $\Phi$. As mentioned above, we regularize the neural attribution by constraining the identified neurons to a subset $\Phi$, in order to prevent degradation in clean response quality after pruning. Given the preferential attribution loss (8), we next detail on the derivation of $\Phi$.
In (8), we can find that the attribution score (3) can be decomposed into,
@@
$a_{t,p}^{i}$ and $a_{t,c}^{i}$ are the scores for poisoned and clean contributions. Correspondinglly, we can have $a_{p}^{i,neu}=max_{t}\,\,a_{t,p}^{i}$ and $a_{c}^{i,neu}=max_{t}\,\,a_{t,c}^{i}$, as neural aggregration in Section 2.2.1. We want to avoid pruning on neurons with significant scores for clean contribution $a_{c}^{i,neu}$, so that the pruned LLM can still generate clean responses that address the user instruction. Since different inputs may come with different magnitudes in scoring, we define normalized scores for clean and poisoned contribution $a_{p}^{i,norm}$ and $a_{c}^{i,norm}$,
$a_{p}^{i,norm}=\frac{a_{p}^{i,neu}}{\sum_{i^{\prime}}a_{p}^{i^{\prime},neu}},\,a_{c}^{i,norm}=\frac{a_{c}^{i,neu}}{\sum_{i^{\prime}}a_{c}^{i^{\prime},neu}}$ (10)
Ideally, we only prune on neurons whose normalized poisoned contribution $a_{p}^{i,norm}$ is much larger than its normalized clean contribution $a_{c}^{i,norm}$. To achieve this, we can define $\Phi$ as,
@@
For example, when $a_{p}^{i,norm}-a_{c}^{i,norm}$ is a large positive value, we can avoid pruning on neurons whose $a_{c}^{i,norm}$ is also large. The importance of pruning with $\Phi$ is demonstrated in Table 5.
@@
Figure 3: The rank of predicted response tokens. Taking (a) as an example, "clean greedy" means the response from greedy decoding is clean. Therefore, the tokens from $y^{*}_{c}$ are always ranked zero. In this case, a poisoned response can be triggered with only one or two tokens. Note that we assume $y_{x}^{p,*}$ starts with answering the injected instructions. We do not find such effect when the injected instructions are answer in the end.